Suppose you deposit $1000 earned from your summer job in a savings account, paying 4.8% compounded annually. How much will you have in your account after 3 years? Write an equation to model the problem and then solve your equation for the given value.
Answer by rfer(16322)

A=P(1+r)^t
A=1000(1+.048)^3
A=1000(1.151)
A=$1151